Performance Products Corporation makes two products, titanium Rims and Posts. Data regarding the two products follow:

   

Direct
Labor-Hours
per unit
Annual
Production
Rims 0.50 18,000 units
Posts 0.60 77,000 units

Additional information about the company follows:

  1. Rims require $12 in direct materials per unit, and Posts require $8.
  2. The direct labor wage rate is $15 per hour.
  3. Rims are more complex to manufacture than Posts and they require special equipment.
  4. The ABC system has the following activity cost pools:
Estimated Activity
Activity Cost Pool Activity Measure Estimated
Overhead
Cost
Rims Posts Total
Machine setups Number of setups $ 34,650 120 80 200
Special processing Machine-hours $ 124,740 2,000 0 2,000
General factory Direct labor-hours $ 648,000 8,000 28,000 36,000

Required:

1. Compute the activity rate for each activity cost pool. (Round your answers to 2 decimal places.)
Answer #1

Machine setup = $34650/200 = $173.25 per nk. Of setup

Special processing = 124740/2000 = 62.36 per machine hour

General factory = 648000 / 36000 = 18 per direct labor hour